OpenStudy (legitemal):

I will Give A Medal To best answer. At a race track, a car of mass 2000 kg crashes into a concrete wall at a speed of 77 m/s. a. If the car comes to a stop when it hits the wall, what is the magnitude of the impulse applied to the car? b. To make things safer, the race track installs barrels of sand along the wall. If the car crashes into the barrels, the time it takes to stop will be increased by a factor of 3, as compared with hitting a concrete wall. How does this affect the force applied to the car during a crash?

OpenStudy (shamim):

a. U know impulse on car=J=F*t=?

OpenStudy (shamim):

U know F*t=mv-mu=?

OpenStudy (shamim):

Here m=mass of ur car v=final velocity of car u=initial velocity of car

OpenStudy (shamim):

m,u,v given in ur question J=?

OpenStudy (anonymous):

But we don't know time ?